# EXCOGITO, an extensible coarse-graining toolbox for the investigation of biomolecules by means of low-resolution representations

### v1

@@ Abstract
Bottom-up coarse-grained (CG) models proved to be essential to complement and sometimes even replace all-atom representations of soft matter systems and biological macromolecules. The development of low-resolution models takes the moves from the reduction of the degrees of freedom employed, that is, the definition of a mapping between a system’s high-resolution description and its simplified counterpart. Even in absence of an explicit parametrisation and simulation of a CG model, the observation of the atomistic system in simpler terms can be informative: this idea is leveraged by the mapping entropy, a measure of the information loss inherent to the process of coarsening. Mapping entropy lies at the heart of the extensible coarse-graining toolbox, or EXCOGITO, developed to perform a number of operations and analyses on molecular systems pivoting around the properties of mappings. EXCOGITO can process an all-atom trajectory to compute the mapping entropy, identify the mapping that minimizes it, and establish quantitative relations between a low-resolution representation and geometrical, structural, and energetic features of the system. Here, the software, which is available free of charge under an open source licence, is presented and showcased to introduce potential users to its capabilities and usage.
@@ 1 Introduction
In the context of soft matter modelling, coarse-graining (CGing) is a broad term encompassing a number of approaches, techniques, and algorithms aimed at constructing low-resolution models of molecular systems 1, 2, 3, 4, 5. The objects of study can range from – structurally – simple molecules (most notably water 6, 7, 8) to very complex biological machineries (proteins, DNA, lipid membranes 9, 10) up to entire cells 11, 12, 13, 14. Such models are conceived so as to entail the necessary amount of information and detail to reproduce specific target properties, and enable the investigation of emergent processes and phenomena on length and time scales that would be out of reach by means of more refined descriptions 15, such as those employing all-atom force fields or even more accurate ab initio approaches.
@@
Mappings can be useful to characterise the system even in absence of a sampling of its conformational space. Indeed, a measure of distance between mappings can be leveraged to highlight structural features of a molecule and explore the mapping space in a quantitative manner 34.
Key to the fruitful usage of these concepts and methods to all applications illustrated insofar, as well as many others, is their implementation in an efficient and easy-to-use software platform. In this work we present, describe, and showcase the extensible coarse-graining toolbox, or EXCOGITO, that was developed to provide users with the necessary instruments to make the most of the concept of mapping and mapping entropy. EXCOGITO implements several tools that allow the investigation of complex molecular systems through various instruments all pivoting on the concept of mapping, leveraging the idea that a relation exists between the most informative low-resolution representation of a system and its physical properties.
In the following we provide the theoretical foundations of the methods implemented in the software, then proceed to describe it and how it works through its application to specific case studies. We will review the previous literature about mapping entropy minimization and mapping space exploration that leverages tools implemented in EXCOGITO, and show the application of these methods to a simple yet nontrivial system, icosalanine.
@@ 2 Methodology
The aim of the EXCOGITO software is to provide users with a simple and efficient tool to investigate the features of the mapping space of a macromolecular system, most notably proteins, and explore the relation they entertain with the physical and biological properties of the latter. This approach relies on a simple yet powerful idea, namely that by losing resolution, e.g. by looking at the MD trajectory in terms of fewer atoms that the total, we gain information about the processes that take place in the system at a more global scale. Since a few years, this idea is put forward by various authors 26, 30, 35, 31, 34, 33, 36; here we do not aim to review the results obtained insofar, nor to discuss the theoretical and practical aspects of its implementation. Rather, we illustrate the software we developed to put this idea at work.
In the following, we will briefly recall the definition and properties of the fundamental theoretical ingredient that lies at the core of EXCOGITO, namely the mapping entropy $S_{map}$, which quantifies the loss of statistical information generated by coarse-graining a macromolecular system. We will then describe the $S_{map}$-based tools that are currently implemented in the software, whose applications range from computing the information loss associated with a specific CG representation to the minimization of $S_{map}$ in the space of possible low-resolution descriptions. For this latter protocol, we will further discuss how a statistical analysis performed over those representations that minimize the information loss with respect to the high-resolution reference can provide insight into the system’s biological properties. Before introducing the concept of mapping entropy and the related EXCOGITO analysis tools, however, it is crucial to understand why and how an information loss arises if one blurs the description of the system of interest. We thus start summarising the basic principles underlying coarse-graining procedures.
Figure 1: Scheme of the EXCOGITO methods, inputs and outputs. The starting point is a MD trajectory, or a given structure thereof; adding the potential energy or the probability associated to each frame one can compute the value of the mapping entropy associated to a given mapping, generate a number of random mappings, and eventually optimise $S_{map}$ through simulated annealing. Alternatively, the trajectory can be analysed through the computation of the norm of a given mapping throughout the run, and the scalar product of two or more mappings can be computed on a given frame.
@@ 2.1 Information loss in coarse-graining
We consider a macromolecular system composed of $n$ constituent atoms—comprising the solute(s) as well as the solvent ones—modeled as point-like particles mutually interacting via classical potentials. The set of Euclidean coordinates $\mathbf{r}_{i}$, i=1,,ni=1,...,n of all atoms defines the high-resolution configuration $\mathbf{r}$, or microstate, of the system to which we associate an atomistic probability distribution $p(\mathbf{r})$. We stress that $p(\mathbf{r})$ is in principle arbitrary, albeit for investigating equilibrium properties one can identify it with, e.g. the canonical distribution, which is
@@
and is thus obtained by integrating over all the high-resolution configurations $\bf r$} of the system that, upon the projection ${\bf M}({\bf r})$, are mapped onto macrostate $\bf R$}, each configuration being weighted with its atomistic probability $p({\bf r})$. Apart from some general features inherited from the fundamental symmetries characterizing the high-resolution system—such as rotational and translational invariance—the CG probability $P(\mathbf{R})$ will critically depend on the prescription employed to group together microscopic configurations to form the macrostates, that is, on the decimation mapping operator $\mathbf{M}$; it is thus easy to imagine that the choice of the mapping will determine how much information on the system properties will be transferred from the high- to the low-resolution representation, and that understanding why a mapping is “better” than another can lead to a deeper understanding of the system.
In the next section we thus address the question of how to select mappings that are “better” than others, specifically starting from the problem of quantifying in an unambiguous manner the quality of a mapping.
@@ 2.2 Mapping entropy and related EXCOGITO tools
In principle, given a CG mapping the associated $S_{map}$ can be directly evaluated through the definition in Eq. 12 provided that the all-atom probability distribution $p(\mathbf{r})$ is known, its backmapped counterpart $\bar{p}_{r}(\mathbf{r)}$ can be explicitly determined and the summation over the microstates $\mathbf{r}$ exhaustively performed. This is the case, for example, of the coarse-graining of simple systems characterized by a finite and low-dimensional configurational space such as discrete classical spins on a small lattice.33 When considering complex macromolecules such as proteins—that is, the main target of the current release of EXCOGITO—however, the calculation of Eq. 12 would require solving analytically intractable high-dimensional integrals over the coordinates of the constituent atoms. Two possible ways of tackling the problem are currently available in EXCOGITO; the associated tools, respectively called measure and measure_kl, enable the user to approximately estimate the mapping entropy associated with the resolution reduction of a macromolecular system starting from a set of ingredients that have to be provided in input to the software.
The first EXCOGITO $S_{map}$ calculation protocol, implemented in the measure tool, relies on configurational sampling and applies to an equilibrium condition in which the high-resolution probability $p(\mathbf{r})$ of the macromolecule is given by the Boltzmann distribution, see Eq. 1. To illustrate the underlying method, let us first consider the case of an arbitrary $p(\mathbf{r})$—whose analytical form is known—and assume that a discrete series of atomistic configurations $\mathbf{r}_{i}$, $i=1,...,K\gg 1$ of the system sampled from such distribution via, e.g., Molecular Dynamics or Monte Carlo simulations is available. Given these configurations and the choice of the CG representation, $S_{map}$ could in principle be estimated as
@@
Eq. 14 shows that $\tilde{S}_{map}$ can be calculated by first computing, for each CG macrostate $\mathbf{R}$, the variance of the atomistic energies of all microscopic configurations that map onto it—the term $\langle(u-\langle u\rangle_{{\bf R}})^{2}\rangle_{{\bf R}}$, where $\langle\cdot\rangle_{{\bf R}}$ is an equilibrium average conditioned to the macrostate. Subsequently, such variances have to be averaged over all possible CG macrostates, each one weighted with its own low-resolution probability $P(\mathbf{R})$. The measure tool of EXCOGITO relies on the estimator in Eq. 14 to evaluate the mapping entropy of a macromolecular system at equilibrium, where a set of high-resolution configuration $\mathbf{r}_{i}$ sampled from the Boltzmann distribution as well as the associated atomistic energies $u(\mathbf{r}_{i})$ have to be provided by the user as input to the software, together with the selected CG representation. We stress that the identification of the CG macrostates $\bf R$} in Eq. 14 is a challenging task: in fact, it cannot be obtained by analytically marginalizing over the discarded degrees of freedom; nor would it be efficient to carry out a restrained sampling where the preserved atoms are kept fixed, since this operation would have to be repeated for a statistically significant number of CG configurations. To circumvent this limitation, measure makes use of a clustering algorithm that, given the available set of high-resolution configurations $\mathbf{r}_{i}$, lumps them in groups based only on the atoms that are retained in the CG mapping, where such groups are then identified with the CG macrostates—a technical analysis of the different prescriptions employed by EXCOGITO to carry out this procedure being reported in Sec. 6.2. Starting from this partitioning, measure computes the energy variance of each macrostate and combines together the results to calculate, via a discretized version of Eq. 14, the mapping entropy associated with a specific CG representation of the system.
In addition to the previously described protocol that is applicable in the case of equilibrium conditions, EXCOGITO further features a different method, implemented in the measure_kl tool, to determine the mapping entropy of a macromolecule. In this second case, a discrete set of high-resolution configurations $\mathbf{r}_{i}$, i=1,,Li=1,...,L of the system as well as the associated probabilities $p(\mathbf{r}_{i})$, with $\sum_{i=1}^{L}p(\mathbf{r}_{i})=1$, have to be provided in input to the software, together with the CG representation chosen. The mapping entropy is then evaluated in measure_kl as a KL divergence over this countable state space, that is,
@@
where the $\wedge$” superscript has been introduced to discriminate this mapping entropy estimator with the cumulant expansion $\tilde{S}_{map}$ one reported in Eq. 14. In Eq. 15, the backmapped probabilities $\bar{p}(\mathbf{r}_{i})$, i=1,,Li=1,...,L are determined, given the selection of the atoms to be retained at the low-resolution level, by clustering all the atomistic configurations into a set of CG macrostates in analogy with the equilibrium framework, see Sec. 6.2 for all technical details. Then, the weight $\bar{p}(\mathbf{r}_{i})$ of the $i$-th configuration is given by the average probability of all microstates that belong to the CG cluster that contains $\mathbf{r}_{i}$.
@@
To provide an example of when this second method can be applied and how the associated ingredients can be obtained, consider a scenario in which, although a series of $K\gg L$ all-atom configurations of the system is available, these represent samples of an unknown distribution $p(\mathbf{r})$. In this case, neither the general estimator in Eq. 13—irrespective of its previously discussed limitations—nor the approximated $\tilde{S}_{map}$ one in Eq. 14 lying at the core of measure can be straightforwardly employed, as they are valid only in equilibrium conditions. One can however perform an atomistic clustering on this ensemble of microstates and lump them in $L$ groups based on similarity criteria. From this, each representative all-atom configuration $\mathbf{r}_{i}$, i=1,,Li=1,...,L in Eq. 15 can then be identified, e.g., with the centroid of a cluster, and the associated probability $p(\mathbf{r}_{i})$ estimated as the fraction of configurations belonging to the cluster. Given these ingredients, the mapping entropy can finally be evaluated through Eq. 13 via an additional (this time, coarse-grained) clustering carried out on the set of $\mathbf{r}_{i}$ starting from a choice of the low-resolution representation of the macromolecule. We underline that measure_kl can also be employed when the functional form of the probability $p(\mathbf{r})$ is known; critically, in the case of equilibrium systems the resulting mapping entropies have been found to correlate with those obtained from the $\tilde{S}_{map}$ estimator in Eq. 14,33 thus highlighting the robustness of this second method.
@@
Given a series of high-resolution configurations of the system of interest endowed with either their all-atom potential energies or their “frequentistic” probabilities, the previously described mapping entropy calculation tools measure and measure_kl enable the user of EXCOGITO to quantify the loss of statistical information experienced by a macromolecule as a consequence of a specific decimation of its microscopic degrees of freedom. One natural requirement would then be that of gauging the quality of such CG representation based on the resulting mapping entropy; at the same time, except for its lower bound, no additional reference $S_{map}$ value can be a priori inferred for an arbitrary system, thus hampering a straightforward interpretation of an information loss calculation performed on a single CG mapping. This problem is further compounded with the dependence of $S_{map}$ on the amount of sites $N$ employed in the simplified description31—the previously introduced degree of coarse-graining, see Eqs. 3 and 7—in addition to their location throughout the molecular structure. For each analyzed system and inspected number of retained sites $N$, it would thus be desirable to identify a “characteristic scale” of $S_{map}$ associated with the somewhat “typical” reduced representations that can be constructed at that degree of coarse-graining; the quality of the proposed CG mapping can then be quantified in terms of the relative information gain/loss that the high-resolution description such selection of sites guarantees, compared to the ones that were chosen as a reference. Critically, in the absence of any previous chemical intuition on the system, the characteristic scale should be as impartial as possible, and it is thus reasonable to deduce it from a totally unbiased exploration of the macromolecule’s mapping space in which low-resolution representations with the desired number of retained sites are randomly probed.
@@
This assessment of the typical spectrum of information loss generated by coarse-graining a macromolecular system can be performed in EXCOGITO through the random or random_kl tools, which respectively rely on the previously discussed mapping entropy estimators $\tilde{S}_{map}$ and $\hat{S}_{map}$ in Eqs. 14, and 13. By providing as input to the software the set of ingredients necessary for a single calculation of $S_{map}$ via the two latter methods, they generate a sequence of CG representations of the system at a fixed degree of coarse-graining in which the $N$ sites are randomly displaced throughout the molecular structure, evaluating the associated mapping entropies. The results of this analysis can then be histogrammed along the $S_{map}$ axis, and the characteristic scale of information loss at the desired value of $N$ can be extracted, e.g., from the average mapping entropy $\mu_{N}$ and standard deviation $\sigma_{N}$ of the sample. Finally, the quality of the proposed CG representation $\mathbf{M}_{N}$ can be gauged in terms of its relative information gain/loss with respect to what would on average arise at the same degree of coarse-graining by randomly choosing the location of the sites; as an example, in Ref. 31 the chosen quantitative quality measure was the standard score $Z(\mathbf{M}_{N})$ of the optimal mapping with respect to the distribution of randomly extracted ones, with
$Z(\mathbf{M}_{N})=\frac{S_{map}(\mathbf{M}_{N})-\mu_{N}}{\sigma_{N}},$ (16)
@@
where we underline that all quantities appearing in Eq. 16 should be calculated always via the same $S_{map}$ estimator in a consistent manner, that is, either through a combination of measure and random to respectively determine $S_{map}(\mathbf{M}_{N})$ and $(\mu_{N},\sigma_{N})$, or alternatively through a combination of measure_kl and random_kl.
@@
With these instruments at hand, we now have all the necessary ingredients to introduce the last $S_{map}$-based analysis tools implemented in EXCOGITO, namely the ones devoted to the identification of the CG representations of the system of interest that, despite a reduction in resolution, are capable of preserving the largest amount of information about the statistical properties of the atomistic reference. Hence, among all the possible decimation mappings that can be designed for a macromolecule at a fixed degree of coarse-graining $N$, we are now looking for those that minimize $S_{map}$; crucially, as it will be discussed in the following, these maximally informative CG representations appear to be related to the system’s functional regions, suggesting the mapping entropy optimization workflow as a promising approach to extract relevant macroscopic insight from raw, all-atom simulation data.

### v2

@@ Abstract
Bottom-up coarse-grained (CG) models proved to be essential to complement and sometimes even replace all-atom representations of soft matter systems and biological macromolecules. The development of low-resolution models takes the moves from the reduction of the degrees of freedom employed, that is, the definition of a mapping between a system’s high-resolution description and its simplified counterpart. Even in the absence of an explicit parametrisation and simulation of a CG model, the observation of the atomistic system in simpler terms can be informative: this idea is leveraged by the mapping entropy, a measure of the information loss inherent to the process of coarsening. Mapping entropy lies at the heart of the extensible coarse-graining toolbox, or EXCOGITO, developed to perform a number of operations and analyses on molecular systems pivoting around the properties of mappings. EXCOGITO can process an all-atom trajectory to compute the mapping entropy, identify the mapping that minimizes it, and establish quantitative relations between a low-resolution representation and the geometrical, structural, and energetic features of the system. Here, the software, which is available free of charge under an open-source licence, is presented and showcased to introduce potential users to its capabilities and usage.
Published on the J. Chem. Inf. Model. on June 11, 2024.
DOI: https://doi.org/10.1021/acs.jcim.4c00490.
@@ 1 Introduction
In the context of soft matter modelling, coarse-graining (CGing) is a broad term encompassing a number of approaches, techniques, and algorithms aimed at constructing low-resolution models of molecular systems 1, 2, 3, 4, 5. The objects of study can range from—structurally—simple molecules (most notably water 6, 7, 8) to very complex biological machineries (proteins, DNA, lipid membranes 9, 10) up to entire cells 11, 12, 13, 14. Such models are conceived so as to entail the necessary amount of information and detail to reproduce specific target properties, and enable the investigation of emergent processes and phenomena on length and time scales that would be out of reach by means of more refined descriptions 15, such as those employing all-atom force fields or even more accurate ab initio approaches.
@@
Mappings can be useful to characterise the system even in the absence of a sampling of its conformational space. Indeed, a measure of distance between mappings can be leveraged to highlight structural features of a molecule and explore the mapping space in a quantitative manner 34.
Key to the fruitful usage of these concepts and methods to all applications illustrated insofar, as well as many others, is their implementation in an efficient and easy-to-use software platform. In this work, we present, describe, and showcase the extensible coarse-graining toolbox, or EXCOGITO, developed to provide users with the necessary instruments to make the most of the concept of mapping and mapping entropy. EXCOGITO, whose workflow and purpose are schematically represented in Fig. 1, implements several tools that allow the investigation of complex molecular systems through various instruments all pivoting on the concept of mapping, leveraging the idea that a relation exists between the most informative low-resolution representation of a system and its physical properties.
Figure 1: A pictorial representation of EXCOGITO and its purpose. The software carries out various types of analyses on a sample of high-resolution, or all-atom, protein configurations. These are investigated in terms of reduced representations, or mappings, in which only a subset of the molecule’s atoms are retained. This kind of analysis can provide novel and useful information on various properties of the system, its conformational variability, its energetics, and how all these properties can be related to the biological function. For further details and examples vide infra as well as Refs. 31, 34.
In the following we provide the theoretical foundations of the methods implemented in the software, then proceed to describe it and how it works through its application to specific case studies. We review the previous literature about mapping entropy minimization and mapping space exploration that leverages tools implemented in EXCOGITO, and show the application of these methods to a simple yet nontrivial system, icosalanine.
@@ 2 Methodology
The aim of the EXCOGITO software is to provide users with a simple and efficient tool to investigate the features of the mapping space of a macromolecular system, most notably proteins, and explore the relation they entertain with the physical and biological properties of the latter. This approach relies on a simple yet powerful idea, namely that by losing resolution, e.g. by looking at the system in terms of fewer atoms than the total, we gain information about the processes that take place at a more global scale. Since a few years, this idea is put forward by various authors 26, 30, 35, 31, 34, 33, 36; here we do not aim to review the results obtained insofar, nor to discuss the theoretical and practical aspects of its implementation. Rather, we illustrate the software we developed to put this idea to work.
In the following, we will briefly recall the definition and properties of the fundamental theoretical ingredient that lies at the core of EXCOGITO, namely the mapping entropy $S_{map}$ quantifying the loss of statistical information generated by coarse-graining a macromolecular system. We will then describe the $S_{map}$-based tools that are currently implemented in the software, whose applications range from computing the information loss associated with a specific CG representation to the minimization of $S_{map}$ in the space of possible low-resolution descriptions. For this latter protocol, we will further discuss how a statistical analysis performed over those representations that minimize the information loss with respect to the high-resolution reference can provide insight into the system’s biological properties. Before introducing the concept of mapping entropy and the related EXCOGITO analysis tools, however, it is crucial to understand why and how an information loss arises if one blurs the description of the system of interest. We thus start summarising the basic principles underlying coarse-graining procedures.
@@ 2.1 Information loss in coarse-graining
We consider a macromolecular system composed of $n$ constituent atoms—comprising the solute(s) as well as the solvent ones—modelled as point-like particles mutually interacting via classical potentials. The set of Euclidean coordinates $\mathbf{r}_{i}$, i=1,,ni=1,...,n of all atoms defines the high-resolution configuration $\mathbf{r}$, or microstate, of the system to which we associate an atomistic probability distribution $p(\mathbf{r})$. We stress that $p(\mathbf{r})$ is in principle arbitrary, in the sense that no assumption needs to be made on the functional form of this probability and/or the nature of the underlying generative process. As a general but relevant example, under the hypothesis of thermal equilibrium one can assume that $p(\mathbf{r})$ is a Boltzmann probability density, i.e. the system is described by the canonical ensemble, with
@@
and is thus obtained by integrating over all the high-resolution configurations $\bf r$} of the system that, upon the projection ${\bf M}({\bf r})$, are mapped onto macrostate $\bf R$}, each configuration being weighted with its atomistic probability $p({\bf r})$. Apart from some general features inherited from the fundamental symmetries characterizing the high-resolution system—such as, e.g., rotational and translational invariance—the CG probability $P(\mathbf{R})$ will critically depend on the prescription employed to group together microscopic configurations to form the macrostates, that is, on the decimation mapping operator $\mathbf{M}$; it is thus easy to imagine that the choice of the CG mapping will determine how much information on the system properties will be transferred from the high- to the low-resolution representation, and that understanding why a mapping is “better” than another can lead to a deeper understanding of the system.
In the next section, we thus address the question of how to select mappings that are “better” than others, specifically starting from the problem of quantifying in an unambiguous manner the quality of a mapping.
@@ 2.2 Mapping entropy and related EXCOGITO tools
In principle, given a CG mapping the associated $S_{map}$ can be directly evaluated through the definition in Eq. 12 provided that the all-atom probability distribution $p(\mathbf{r})$ is known, its backmapped counterpart $\bar{p}_{r}(\mathbf{r)}$ can be explicitly determined and the summation over the microstates $\mathbf{r}$ exhaustively performed. This is the case, for example, of the coarse-graining of simple systems characterized by a finite and low-dimensional configurational space such as discrete classical spins on a small lattice 33. When considering complex macromolecules such as proteins—that is, the main target of the current release of EXCOGITO—however, even in the case of a known functional form of the probability distribution, the calculation of Eq. 12 would require solving analytically intractable high-dimensional integrals over the coordinates of the constituent atoms.
Two possible ways of tackling the problem are currently available in EXCOGITO; the associated tools, respectively called measure and measure_kl, enable the user to approximately estimate the mapping entropy associated with the resolution reduction of a macromolecular system starting from a set of ingredients that have to be provided in input to the software. As we illustrate hereafter, these ingredients amount to the configurations of the system under examination, and either the potential energy associated to each configuration (measure tool), or the value of its statistical weight (measure_kl tool). In the following subsections, we describe these different approaches, their conceptual frameworks, and contexts of applicability.
CUMULANT EXPANSION STRATEGY — The first EXCOGITO $S_{map}$ calculation protocol, implemented in the measure tool, relies on configurational sampling and applies to an equilibrium condition in which the high-resolution probability $p(\mathbf{r})$ of the macromolecule is given by the Boltzmann distribution, see Eq. 1. To illustrate the underlying method, let us first consider the case of an arbitrary $p(\mathbf{r})$—whose analytical form is known—and assume that a discrete series of atomistic configurations $\mathbf{r}_{i}$, $i=1,...,K\gg 1$ of the system sampled from such distribution via, e.g., Molecular Dynamics or Monte Carlo simulations is available. Given these configurations and the choice of the CG representation, $S_{map}$ could in principle be estimated as
@@
Eq. 14 shows that $\tilde{S}_{map}$ can be calculated by first computing, for each CG macrostate $\mathbf{R}$, the variance of the atomistic energies of all microscopic configurations that map onto it—that is to say, the term $\langle(u-\langle u\rangle_{{\bf R}})^{2}\rangle_{{\bf R}}$, where $\langle\cdot\rangle_{{\bf R}}$ is an equilibrium average conditioned to the macrostate. Subsequently, such variances have to be averaged over all possible CG macrostates, each one weighted with its own low-resolution probability $P(\mathbf{R})$. The measure tool of EXCOGITO relies on the estimator in Eq. 14 to evaluate the mapping entropy of a macromolecular system at equilibrium, where a set of high-resolution configuration $\mathbf{r}_{i}$ sampled from the Boltzmann distribution as well as the associated atomistic energies $u(\mathbf{r}_{i})$ have to be provided by the user as input to the software, together with the selected CG representation. We stress that the identification of the CG macrostates $\bf R$} in Eq. 14 is a challenging task: in fact, it cannot be obtained by analytically marginalizing over the discarded degrees of freedom; nor would it be efficient to carry out a restrained sampling where the preserved atoms are kept fixed, since this operation would have to be repeated for a statistically significant number of CG configurations. To circumvent this limitation, measure makes use of a clustering algorithm that, given the available set of high-resolution configurations $\mathbf{r}_{i}$, lumps them in groups based only on the atoms that are retained in the CG mapping, where such groups are then identified with the CG macrostates—a technical analysis of the different prescriptions employed by EXCOGITO to carry out this procedure being reported in Sec. 6.2. Starting from this partitioning, measure computes the energy variance of each macrostate and combines together the results to calculate, via a discretized version of Eq. 14, the mapping entropy associated with a specific CG representation of the system.
KULLBACK-LEIBLER STRATEGY — In addition to the previously described protocol that is applicable in the case of equilibrium conditions, EXCOGITO further features a different method, implemented in the measure_kl tool, to determine the mapping entropy of a macromolecule. In this second case, a discrete set of high-resolution configurations $\mathbf{r}_{i}$, i=1,,Li=1,...,L of the system as well as the associated probabilities $p(\mathbf{r}_{i})$, with $\sum_{i=1}^{L}p(\mathbf{r}_{i})=1$, have to be provided in input to the software, together with the CG representation chosen. The mapping entropy is then evaluated in measure_kl as a KL divergence over this countable state space, that is,
@@
where the hat superscript has been introduced to discriminate this mapping entropy estimator with the cumulant expansion $\tilde{S}_{map}$ equilibrium one reported in Eq. 14. In Eq. 15, the backmapped probabilities $\bar{p}(\mathbf{r}_{i})$, i=1,,Li=1,...,L are determined, given the selection of the atoms to be retained at the low-resolution level, by clustering all the atomistic configurations into a set of CG macrostates in analogy with the equilibrium framework, see Sec. 6.2 for all technical details. Then, the weight $\bar{p}(\mathbf{r}_{i})$ of the $i$-th configuration is given by the average probability of all microstates that belong to the CG cluster that contains $\mathbf{r}_{i}$.
@@
To provide an example of when this second method can be applied and how the associated ingredients can be obtained, consider a scenario in which, although a series of $K\gg L$ all-atom configurations of the system is available, these represent samples of an unknown distribution $p(\mathbf{r})$—e.g., they are obtained from experimental measurements or emerge from MD simulations of a system in a stationary but non-equilibrium condition. In this context, neither the general estimator in Eq. 13 (irrespective of its previously discussed limitations) nor the approximated $\tilde{S}_{map}$ one in Eq. 14 lying at the core of measure can be straightforwardly employed. However, one can perform an atomistic clustering on this ensemble of microstates and lump them in $L$ groups based on similarity criteria. From this, each representative all-atom configuration $\mathbf{r}_{i}$, i=1,,Li=1,...,L in Eq. 15 can then be identified, e.g., with the centroid of a cluster, and the associated probability $p(\mathbf{r}_{i})$ estimated as the fraction of configurations belonging to the cluster. Given these ingredients, the mapping entropy can finally be evaluated through Eq. 13 via an additional (this time, coarse-grained) clustering carried out on the set of $\mathbf{r}_{i}$ starting from a choice of the low-resolution representation of the macromolecule. We underline that measure_kl can also be employed when the functional form of the probability $p(\mathbf{r})$ is known; critically, in the case of equilibrium systems the resulting mapping entropies have been found to correlate with those obtained from the $\tilde{S}_{map}$ estimator in Eq. 14,33 thus highlighting the robustness of this second method.
@@
ESTIMATION OF THE QUALITY OF A CG MAPPING — Given a series of high-resolution configurations of the system of interest endowed with either their all-atom potential energies or their “frequentistic” probabilities, the previously described mapping entropy calculation tools measure and measure_kl enable the user of EXCOGITO to quantify the loss of statistical information experienced by a macromolecule as a consequence of a specific decimation of its microscopic degrees of freedom. One natural requirement would then be that of gauging the quality of such CG representation based on the resulting mapping entropy; at the same time, except for its lower bound, no additional reference $S_{map}$ value can be a priori inferred for an arbitrary system, thus hampering a straightforward interpretation of an information loss calculation performed on a single CG mapping. This problem is further compounded with the dependence of $S_{map}$ on the amount of sites $N$ employed in the simplified description31—the previously introduced degree of coarse-graining, see Eqs. 3 and 7—in addition to their location throughout the molecular structure. For each analysed system and inspected number of retained sites $N$, it would thus be desirable to identify a “characteristic scale” of $S_{map}$ associated with the somewhat “typical” reduced representations that can be constructed at that degree of coarse-graining; the quality of the proposed CG mapping can then be quantified in terms of the relative information gain/loss that the high-resolution description such selection of sites guarantees, compared to the ones that were chosen as a reference. Critically, in the absence of any previous chemical intuition on the system, the characteristic scale should be as impartial as possible, and it is thus reasonable to deduce it from a totally unbiased exploration of the macromolecule’s mapping space in which low-resolution representations with the desired number of retained sites are randomly probed.
@@
This assessment of the typical spectrum of information loss generated by coarse-graining a macromolecular system can be performed in EXCOGITO through the random or random_kl tools, which respectively rely on the previously discussed mapping entropy estimators $\tilde{S}_{map}$ and $\hat{S}_{map}$ in Eqs. 14 and 15. By providing as input to the software the set of ingredients necessary for a single calculation of $S_{map}$ via the two latter methods, random or random_kl generate a sequence of CG representations of the system at a fixed degree of coarse-graining in which the $N$ sites are randomly displaced throughout the molecular structure, evaluating the associated mapping entropies. The results of this analysis can then be histogrammed along the $S_{map}$ axis, and the characteristic scale of information loss at the desired value of $N$ can be extracted, e.g., from the average mapping entropy $\mu_{N}$ and standard deviation $\sigma_{N}$ of the sample. Finally, the quality of the proposed CG representation $\mathbf{M}_{N}$ can be gauged in terms of its relative information gain/loss with respect to what would on average arise at the same degree of coarse-graining by randomly choosing the location of the sites; as an example, in Ref. 31 the chosen quantitative quality measure was the standard score $Z(\mathbf{M}_{N})$ of the optimal mapping with respect to the distribution of randomly extracted ones, with
$Z(\mathbf{M}_{N})=\frac{S_{map}(\mathbf{M}_{N})-\mu_{N}}{\sigma_{N}}.$ (16)
@@
We underline that all quantities appearing in Eq. 16 should be calculated always via the same $S_{map}$ estimator in a consistent manner, that is, either through a combination of measure and random to respectively determine $S_{map}(\mathbf{M}_{N})$ and $(\mu_{N},\sigma_{N})$, or alternatively through a combination of measure_kl and random_kl.
@@
CG MAPPING OPTIMISATION — With these instruments at hand, we now have all the necessary ingredients to introduce the last $S_{map}$-based analysis tools implemented in EXCOGITO, namely the ones devoted to the identification of the CG representations of the system of interest that, despite a reduction in resolution, are capable of preserving the largest amount of information about the statistical properties of the atomistic reference. Hence, among all the possible decimation mappings that can be designed for a macromolecule at a fixed degree of coarse-graining $N$, we are now looking for those that minimize $S_{map}$; crucially, as it will be discussed in the following, these maximally informative CG representations appear to be related to the system’s functional regions, suggesting the mapping entropy optimization workflow as a promising approach to extract relevant macroscopic insight from raw, all-atom data.